OpenStudy (anonymous):

Simplify sqrt(9x^4)/36

OpenStudy (bloomlocke367):

is this your problem\[\frac{ \sqrt{9x^4} }{36 }\]

OpenStudy (bloomlocke367):

@xoxoLoveBre

OpenStudy (anonymous):

Yes

OpenStudy (bloomlocke367):

okay, do you know what to do first?

OpenStudy (anonymous):

Not really. I'm honestly just super confused about it all

OpenStudy (bloomlocke367):

okay, well can you think of anything you can do to get rid of the radical?

OpenStudy (anonymous):

Wait the 36 is inside the sqrt symbol too. So isn't the whole thing a radical?

OpenStudy (bloomlocke367):

oh, it is? okay. I meant the radical sign. so, since the 36 is under it too, you want to start by breaking it up. Meaning:\[\frac{ \sqrt{9x^4} }{ \sqrt{36} }\]

OpenStudy (anonymous):

Ok well the square root of 36 is 6.

OpenStudy (bloomlocke367):

okay. so now you have\[\frac{ \sqrt{9x^4} }{ 6 }\]

OpenStudy (anonymous):

Ok and the square root of 9 is 3

OpenStudy (bloomlocke367):

okay... what about x^4? x^4 is the same as x*x*x*x? how many groups of 2 can you make with x?

OpenStudy (anonymous):

1 group of 2?

OpenStudy (bloomlocke367):

no... how many groups is that?